OpenStudy (anonymous):

Laila bought 20 strawberries. She ate 1 strawberry. What percentage of the strawberries did Laila eat? 2% 3% 4% 5%

OpenStudy (anonymous):

@JessicaAngel @LynFran @campbell_st

OpenStudy (anonymous):

Think of it as a proportion. X/100 = 1/20

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

@0400782 I no I don't know wat dat is can just give me da answer its a test and I have a minute left plzzz

OpenStudy (lynfran):

what 1/20*100=?

OpenStudy (anonymous):

5

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

thx